Solve
a) x^3 + 8x^2 +17x + 10 = 0
Can you please help me out? Thanks so much in advance:)
Can you also show me the steps it would really help me understand

Answer by stanbon(75887)

Solve
a) x^3 + 8x^2 +17x + 10 = 0
f(-1) = -1 + 8 - 17 + 10 = 0
So x = -1 is a root.
Use synthetic division to find the other roots:
-1)....1....8....17....10
.......1....7....10....|..0
------
Solve the Quotient:
x^2 + 7x + 10 = 0
(x+5)(x+2) = 0
----
All the roots are x = -1,x = -5, x = -2
===========================================
Cheers,
Stan H.
================